# Multi-directional investigations on quiet time suprathermal ions measured by ASPEX-STEPS on-board Aditya L1

### v1

@@ 6 Summary
In this study we focus on short-term quiet periods of suprathermal ions during solar maximum and the analysis leads to the following key findings:
@@
    The distribution of suprathermal ions during quiet periods is isotropic, as indicated by the similar spectral index values ($\sim 2$) observed in all four directions—Parker spiral (PS), intermediate (IM), Earth pointing (EP), and north pointing (NP).
@@
    This isotropy supports the assumption made by L. Fisk & G. Gloeckler (2008), who considered an isotropic suprathermal ion distribution while solving the Parker transport equation to explain the acceleration of suprathermal ions.
@@
    The suprathermal ion pool during quiet periods are predominantly composed of residual particles from previous impulsive and gradual solar energetic particles (SEPs) events in solar maximum conditions.
@@
    The ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ abundance ratio indicates that approximately 25% of the quiet periods exhibit signatures of impulsive SEPs (ISEPs) contributions.
@@
    The $\mathrm{Fe}/\mathrm{O}$ ratio suggests that about 45% of the suprathermal ion population during quiet times can be attributed to ISEPs.
  6. 6.
    The $\mathrm{C}/\mathrm{O}$ ratio reveals that the suprathermal population comprises approximately 20% contribution from ISEPs and about 50% from gradual SEPs (GSEPs) events.
Authors also would like to thank various centres of ISRO for providing technical support and facilities for the test and calibration of ASPEX payload. Thanks are also due to the project, mission, and review teams of ISRO for their support. Authors are also grateful to the principal investigators and members of the ACE-ULEIS team for generating and managing the datasets used in this work. This work is supported by the Department of Space, Government of India. We greatly appreciate the discussion and suggestions given by Mihir Desai from Southwest Research institute (SWRI).
@@
The data used in this study are publicly available. The ASPEX-STEPS data from the Aditya-L1 mission are accessible from two sources: data from January 7 to May 10, 2024 are available via Zenodo (https://zenodo.org/records/15833452), and data from May 11, 2024 onwards can be obtained from the ISSDC PRADAN portal (https://pradan.issdc.gov.in/al1/). The ULEIS data from the ACE spacecraft are available through NASA’s CDAWeb (https://cdaweb.gsfc.nasa.gov/).

### v2

@@ 6 Summary
In this study we focus on short-term quiet time of suprathermal ions during solar maximum and the analysis leads to the following key findings:
@@
    The distribution of suprathermal ions during quiet time is nearly isotropic, as indicated by the similar spectral index values within the range of $1.84\pm 0.09$ to $2.15\pm 0.02$ (average value $1.99\pm 0.06$) observed in all four directions—Parker spiral (PS), intermediate (IM), Earth pointing (EP), and north pointing (NP).
@@
    This near isotropic behavior supports the assumption made by L. Fisk & G. Gloeckler (2008), who considered an isotropic suprathermal ion distribution while solving the Parker transport equation to explain the acceleration of suprathermal ions.
@@
    We infer that this near isotropic behavior comes out owing to the consideration of species-integrated fluxes, short time-averaging and fixing a year within a solar cycle, These choices minimize the variable factors that may contribute to the quiet time variability of the spectral indices.
@@
    Elemental abundance ratios indicate that the quiet-time suprathermal ion population has significant contributions from solar energetic particle (SEP) events: approximately 25% from impulsive SEPs based on the $^{3}\mathrm{He}/{}^{4}\mathrm{He}$ ratio, about 45% from impulsive SEPs inferred from the $\mathrm{Fe}/\mathrm{O}$ ratio, and nearly 20% from impulsive SEPs along with  50% from gradual SEPs as suggested by the $\mathrm{C}/\mathrm{O}$ ratio.
@@
    The suprathermal ion pool during quiet time are predominantly composed of residual particles from previous impulsive and gradual solar energetic particles (SEPs) events in solar maximum conditions.
Authors also would like to thank various centres of ISRO for providing technical support and facilities for the test and calibration of ASPEX payload. Thanks are also due to the project, mission, and review teams of ISRO for their support. Authors are also grateful to the principal investigators and members of the ACE-ULEIS team for generating and managing the datasets used in this work. We also sincerely acknowledge Dr. Mihir Desai from the Southwest Research Institute (SwRI) for discussions during the course of this work.
@@
The data used in this study are publicly available. The ASPEX-STEPS data from the Aditya-L1 mission are accessible from two sources: data from January 7 to May 10, 2024, are available via Zenodo (10.5281/zenodo.15833452) (D. Chakrabarty, 2025), and data from May 11, 2024, onwards can be obtained from the ISSDC PRADAN portal (https://pradan.issdc.gov.in/al1/). The ULEIS data from the ACE spacecraft are available through NASA’s CDAWeb (https://cdaweb.gsfc.nasa.gov/) under the DOI 10.48322/agmq-ex61 (G. M. Mason et al., 2022).